Suppose babies born after a gestation period of 32 to 35 weeks have a mean weight of 2500 grams and a standard deviation of 700 grams while babies born after a gestation period of 40 weeks have a mean weight of 3000 grams and a standard deviation of 480 grams. If a 33​-week gestation period baby weighs 2650 grams and a 41​-week gestation period baby weighs 3150 ​grams, find the corresponding​ z-scores. Which baby weighs more relative to the gestation​ period?
Answer #1

Z-score corresponds to  a 33​ to 35 week gestation period baby weighs 2650 grams is,

Z = (2650 - 2500)/700 = 0.2143

Z-score = 0.21

and Z-score corresponds to a 41​-week gestation period baby weighs 3150 ​grams is,

Z = (3150 - 3000)/480 = 0.3125

Z-score = 0.31

Here, z = 0.31 > z = 0.21

Therefore, the baby born in week 41 weighs more relative to the gestation period, since its zscore is larger.